how much 2% milk and 9% whipping cream need to be mixed to make 35 gallons of 4% milk?
Answer by stanbon(75887)

how much 2% milk and 9% whipping cream need
to be mixed to make 35 gallons of 4% milk?
----
Equation:
fat + fat = fat
0.02x + 0.09*(35-x) = 0.04(35)
-----
Multiply thru by 100 to get:
2x + 9*35 - 9x = 4*35
-7x = -5*35
---
x = 25 gallons (amt. of 2% milk needed)
---
35-x = 10 gallons (amt. of 9% milk needed)
=====================
cheers,
Stan H.
=====================